• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M.Olson

    Log entry time 08:02:55 on May21,2007

    Entry number 202842

    keyword=Owl Shift Summary

    Owl Shift Summary:

    M. Olson, B. Hahn

    Run 3671 in progress at beginning of shift (He3 Production)
    0027: End run 3671 (1.25M)

    0114: End run 3672 (1.25M)

    0203: End run 3673 (1.25M)
    0258: End run 3674 (1.25M)

    0359: Beam very rocky over last 40-ish minutes. Called MCC and offered to run at 90uA to see if that helps the stability. If it does, we will still benefit overall from the increase in efficiency.

    0406: End run 3675 (1.25M, many beam trips during run)

    0520: End run 3676 (1.25M, significant beam off time during run)

    0624: End run 3677 (1.25M)
    0655: Queried MCC about going back to 105uA if Hall C comes back on at 40uA. If they come back at 80uA, then I feel we are best to stay at 90uA.

    0715: Hall C coming back at 80uA, so we will stay at 90uA for now.
    0721: End run 3678 (1.25M)

    Shift ends on run 3679